OpenStudy (anonymous):

HELP FOR A MEDAL PLSSSS REALLY NEED HELP During a lab, you heat 1.62 g of a CoCl2 hydrate over a Bunsen burner. After heating, the final mass of the dehydrated compound is 0.88 g. Determine the formula of the hydrate and also give the full name of the hydrate. Please help I'm having so much trouble with this question

OpenStudy (anonymous):

determine the gram of h2o : 1.62- o.88 = 0.74 determine the mole of Cocl2: 0.88/ 130= 6.73*10^-3 ( 130 is molar mass) dertermine the mole of h2o :0.74/18=0.0411 determine the ratio : cocl2: 6.73*10^-3/(6.72*10^-3)=1 h2O: 0.041/(6.72*10^-3)=6 so the formula will be CoCl2 6 h2o, we called the cobal chlorate hexan hydrate